OpenStudy (anonymous):

convert r(θ)=2/(2cos(θ)+3sin(θ)) to a cartesian function

myininaya (myininaya):

you have the following equations to use: \[x=rcos(\theta)\] \[y=rsin(\theta)\] \[x^2+y^2=r^2\] \[\tan(\theta)=\frac{y}{x}\]

myininaya (myininaya):

do you might want to try moving the cos sin part with the r part to try to use those first two equation